Can you please help me solve this, the product of two consecutive even integers decreased by 60 is 20. write an equation that represents this situation and solve it to find the integers.
Answer by stanbon(75887)

Can you please help me solve this, the product of two consecutive even integers decreased by 60 is 20. write an equation that represents this situation and solve it to find the integers.
Equation::
(2x)(2x+2)-60 = 20
4x^2 + 4x -80 = 0
x^2 + x - 20 = 0
(x+5)(x-4) = 0
x = -5 or x = 4
----
2x = -10 ; 2x+2 = -8
Or
2x = 8 ; 2x+2 = 10
------------
Cheers,
Stan H.
--------------